OpenStudy (thomas5267):

Need hints on this question. Express \(\sqrt{3}\sin(x)+\cos(x)\) in the form \(R\sin(x+\alpha)\).

OpenStudy (thomas5267):

\[ R\sin(x+\alpha)=R\left(\sin(x)\cos(\alpha)+\sin(\alpha)\cos(x)\right) \] What to do next?

OpenStudy (vishweshshrimali5):

Divide amd multiply each term by 2

OpenStudy (vishweshshrimali5):

*and

OpenStudy (vishweshshrimali5):

I am sure that was what Ganeshi was going to say , most probably :)

ganeshie8 (ganeshie8):

yes :) \[\large \sqrt{3}\sin(x)+\cos(x) = \langle \cos x , \sin x \rangle \bullet \langle 1, \sqrt{3} \rangle \]

OpenStudy (vishweshshrimali5):

See ! ^^^ :D

OpenStudy (vishweshshrimali5):

You will get this on dividing and multiplying by 2: \[\large{2(\cfrac{\sqrt{3}}{2} \sin x + \cfrac{1}{2} \cos x)}\] Notice the sqrt(3)/2 and 1/2

OpenStudy (vishweshshrimali5):

Seem familiar ? @thomas5267

OpenStudy (vishweshshrimali5):

cos(30 deg) = sqrt(3)/2 sin(30 deg) = 1/2

OpenStudy (vishweshshrimali5):

The expression would thus become: \[\large{2(\cos \cfrac{\pi}{6} \sin x + \sin \cfrac{\pi}{6} \cos x)}\] Compare this with : \(\color{blue}{\text{Originally Posted by}}\) @thomas5267 \[ R\sin(x+\alpha)=R\left(\sin(x)\cos(\alpha)+\sin(\alpha)\cos(x)\right) \] What to do next? \(\color{blue}{\text{End of Quote}}\)